Let  be a linear transformation. Let  and  be zero vectors of  and. Show that


1
Expert's answer
2021-09-07T01:25:46-0400

"Let \ T: \mathbb{R}^{n} \rightarrow \mathbb{R}^{m} \\"  be a linear transformation.

"Let \ \mathbf{0}_{n} \ and \ \mathbf{0}_{m} \ be \ zero \ vectors \ of \ \mathbb{R}^{n} \ and \ \mathbb{R}^{m} , respectively. \\Show \ that \ T\left(\mathbf{0}_{n}\right)=\mathbf{0}_{m}."


"Observe \ that \ we \ have\ 0 \cdot \mathbf{0}_{n}=\mathbf{0}_{n} .\\ (This \ is \ a \ scalar \ multiplication \ of \ the \ scalar \ 0 and \ the \ vector \ \mathbf{0}_{n} \\
Now \ we \ have\ 
T\left(\mathbf{0}_{n}\right)=T\left(0 \cdot \mathbf{0}_{n}\right)=0 \cdot T\left(\mathbf{0}_{n}\right)=\mathbf{0}_{m}\\
 
Here \ we \ used \ the \ one\ of\ the \ properties \ of \ the\ linear\ transformation \ T \ \\ in \ the\ second \ equality. \\
Note \ that \ \mathbf{0}_{n}=\mathbf{0}_{n}-\mathbf{0}_{n} \\
Thus \ we \ have \\
 
T\left(\mathbf{0}_{n}\right)=T\left(\mathbf{0}_{n}-\mathbf{0}_{n}\right)=T\left(\mathbf{0}_{n}\right)-T\left(\mathbf{0}_{n}\right)=\mathbf{0}_{m}\\
 
where \ we \ used \ the \ linearity \ of \ T \ in \ the \ second \ equality.\\
In \ the \ last \ equality, \ note \ that \ the \ vector \ T\left(\mathbf{0}_{n}\right) \ is \ m -dimensional \ vector."